OpenStudy (anonymous):

Regard y as the independent variable and x as the dependent variable and use implicit differentiation to find dx/dy. x^3y^2 − x^4y + 4xy^3 = 0

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Ooo they're switching up the x and y's on you. So sneaky! :O

OpenStudy (anonymous):

I know! lol do I just do a normal product rule?

zepdrix (zepdrix):

Yes, start with normal product rule. But since we're taking the derivative `with respect to y`, any `non-y` derivative we take will result in an extra little nugget. So like when you take the derivative of a function of x, you should also get an x' attached to it.

OpenStudy (anonymous):

so like 3x^2dx/dy?

zepdrix (zepdrix):

yes good.

zepdrix (zepdrix):

Primes are better though ^^ Less cumbersome lol.

OpenStudy (anonymous):

3x^2' * y^2 + 2y * x^3 - 4x^3' * y + x^4 * 1 +4' * y^3 + 3y^2 * 4x

zepdrix (zepdrix):

Ok a few boo boos to fix up real quick.

zepdrix (zepdrix):

The prime goes on a separate x, not on the one you differentiated.\[\Large\bf\sf 3x^2x'y^2+2x^3y-4x^3x'y\color{red}{+}x^4(1)+4x'y^3+12xy^2=0\]

zepdrix (zepdrix):

We also need to fix this red part. See how there is a negative in front of your second term? That negative gets `distributed` to each term in your product rule setup.

zepdrix (zepdrix):

\[\Large\bf\sf \frac{dx}{dy}\quad=\quad x'\]

OpenStudy (anonymous):

wait so 3x^2x′y^2 +2x^3 y−4x^3x′y - x^4 (1) - 4x′y^3 -12xy^2 =0 ???

zepdrix (zepdrix):

Mmmmm yah that looks good! :D

OpenStudy (anonymous):

:) yay!!

zepdrix (zepdrix):

And now we need to solve for x'. Any confusion so far? :U

OpenStudy (anonymous):

a little but you explain well so I am ok lol

OpenStudy (anonymous):

alright how do I solve for x' combine like terms?

zepdrix (zepdrix):

You won't be able to combine like terms since there is a big mess of x's and y's. We'll instead need to `factor out` an x' from each term that contains an x'.

OpenStudy (anonymous):